OpenStudy (johnweldon1993):

Alright so let me check If I have this right...Electric Field of Semicircle...

OpenStudy (johnweldon1993):

Find the electric field due to a half circle of radius R and total charge Q. (a) First estimate by dividing the charge into five equal parts. (b) Repeat for 9 equal parts, (c) Using integral calculus solve exactly. So we would have  a) Break into 5 equal pieces

OpenStudy (johnweldon1993):

 Total Charge = Q So charge here is Q/5 Charge per unit length =\(\large \frac{Q}{L}\) Q will be 1/5Q and L would be half the circle divided into 1/5 parts...so... \(\large \frac{1}{5} \frac{(1/5)Q}{\pi r} = \frac{(1/5)Q}{5 \pi r}\)

OpenStudy (johnweldon1993):

Let that = \(\large \lambda\) So \[\large \lambda = \frac{.2Q}{5\pi r}\] Now charge on each slice would be \[\large dQ = \lambda Rd\theta\] So we would have a small portion of the electric field as \[\large dE = \frac{kdQ}{r^2} = \frac{k\lambda }{r}d\theta\]

OpenStudy (johnweldon1993):

Components of each slice are \[\large dE_x = dEcos \theta\] \[\large dE_y = dEsin\theta\] So \(\large E_x\) would be \[\large \int^{\pi}_{0} \frac{k \lambda}{r}cos \theta d\theta = \frac{k \lambda}{r} \int^{\pi}_{0} cos\theta d\theta = \frac{k \lambda}{r}sin\theta \left.\right|_0^\pi\] And \(\large E_y\) would be \[\large \int^{\pi}_{0} \frac{k \lambda}{r}sin \theta d\theta = \frac{k \lambda}{r} \int^{\pi}_{0} sin\theta d\theta = -\frac{k \lambda}{r}cos\theta \left.\right|_0^\pi\] Right? and then just solve knowing \(\large \lambda = \frac{.2Q}{5\pi r}\)

OpenStudy (johnweldon1993):

And for 'b' just replace that 5 on the denominator with 9 instead...and finally integration just using \(\large \pi r\)as the denominator since we want it all?